2012 March 12 UT LBC/MODS/LUCI checkout

Cirrus at sunset, but not too extensive...

LBC

18:50 Bz flats with LBC. z flat has strange pattern, some shadow across center. This was PA=0 20120312. up to 015334. Trying PA=180. Start with 015722... 020059. This has a different gradient but odd. Brightest in upper right (1,4).

19:00 Switch to gi filters and back to PA=0. First i will be 020245... similar pattern as before, but as it gets darker this pattern diminishes. Could be scattered light from the dome or structure? Moon is not yet up.

19:19 Collimate at this flat field position (it is full of stars, 05:24 27deg, and better to choose blank sky). Amb temp is 6.5 deg warmer than 24 hrs ago, but mirrors are close to equilibrium now and first dofpia, /first fit pupils but gave large +z4. Second iteration is good for both sides. Converged in 4-5 iterations. 1.4"/1.1" in B/R seeing.

19:32 WT10_301 pointing correction. lbcrangebal... AO offsets:
side T1 T2 Z RX RY
left -0.96 -0.41 0.715 13.728 -1.483
right -0.96 -0.47 0.497 17.019 -0.322

lbcrangebal:
side dX dY dRX dRY
SX 0.025 0.021 0.442 -0.535
DX -0.025 -0.288 -6.183 0.540

delta CE = 3.931 delta CA = 6.347

19:52 Slew to SA98, dofpia

AO offsets:
side T1 T2 Z RX RY
left -1.01 -0.28 0.568 11.640 -3.206
right -1.01 -0.34 0.368 16.509 -0.130

20:06 Slew to field of RLQ MG2J0624... to test guiding/focuscor/dithering. 3 dither positions, 2 filters (nexp=2) at each.

AO offsets:
side T1 T2 Z RX RY
left -1.07 -0.55 0.651 11.076 -5.521
right -1.07 -0.37 0.356 12.692 0.985

Red channel is drifting out of focus, even though there are corrections being sent from the tech chip pupils. Guide star is visibly out of focus now. Will have to look at sequence of guide images offline. See, e.g. 032959 which shows the "tilt" pattern which we had seen and examined in the fall of 2011 (IT 3875).

Immediately after OB finished run dofpia:

20:37 MST First red and blue images are converged.

20:39 MODS

From LBTO account, we have a problem running MODS.

LBTO@obs2:6 % mods1 status
LBTO@192.168.139.130's password:

LBTO@obs2:7 % mods1 start gui
LBTO@192.168.139.130's password:

(entering the LBTO password does not work).

Called Michele who tried to reach Stephen Hooper. We can run MODS from LBTB's account and so are doing that for the moment.

21:20 modsWake, sieveSnap. irLaser is not putting out any power POUT=0.0.

22:30 Around this time, finally did a graceful shutdown, shutting down the Red, Blue and mods1data computers. I left mods1 computer running. Now, when I started up these computers, I got an error with mods1data not being able to start IE. And an ssh error to the mods1 computer
ssh: connect to host 192.168.139.130. No route to host.

23:00 Going to LUCI

** Try vanBoekel program acquisition. Needed to rotate 0.23deg to align both stars in slit. This offset to be added to PA in acq script.

00:00 deCarli standard star acquisition and observation Y=1068.5 on slit. J and 210zJHK. ~17,000 counts peak, seeing 0.7". But sinking... centered reference star on slit and the acq script made the blind offset. Going to take a pair of exposures 60x4 since we are here and it is a check to see whether the spectrum appears, although the airmass is just at the limit (<1.5).

01:15 Continuing with deCarli script. Elevation is low, but seeing is good. Stopped after 10 of 20 offsets - this was just to give us practice and to give the PI a chance to check the data - is the object spectrum there?

01:31 Now going to take calib/arc at position. Need to use web io to put in calib unit. No, it is IN. But there are error messages in the log which refer to Errors:
Error receiving data from WEB_IO, socket connection on 192.168.0.12, port 80.
Error setting the WebIO output channel 2 to true: 192.168.0.12: rmi://localhost:1098/WebIO_2.

Halo2: 10sec DIT gives 1700 counts. Not much. Try halo1: 10sec DIT gives 6000 counts.

Edited scriptA_flat to use Halo1 and 20sec DIT. Ran this, need to run one more time because I forgot to turn off lamp manually and did not get the leading dark.

But the leading dark was done and was 75 - 73/74, 75/76 and now 77/78. 75 has a strange pattern in the dark, a steep gradient towards the outer regions.

02:00 Slewing to vanBoekel field again. Ran acquisition, with new PA the two stars are well aligned. Centered and started/exited science script. Started/exited sky exposure. OK...

02:34 Try MPE field. Took slit image (88) and then wanted to offset 10",10" to take a sky field image. But the guiding failed because the probe failed to reach position. David is restarting the AGw.

02:50 Another slit image. Now the offset completed. But the -10,-10 offset did not complete without error.

Finally worked - with 0,5 and 0,-5 offsets. But the faint galaxy could not be seen in DIT=10, NDIT=2, so we increased NDIT to 15. This pair will be 94 (sky) and 95 (onsrc). Found that we needed to move 1.32" to the right (+X along the detector coord). 96 is the confirmatory image with the mask still in turnout. Slit is at 1022 (top) to 1018 (bottom). The star is at 1010.5, 1433.4 and the galaxy at 1008,753.

03:54/10:54 UT between image 99 and 100 guide probe failed

13:20 UT Finished 20 dithers on the MPE field. The guide probe failed on 4 positions. Concluded with telluric, flat and arc.

13:14 UT Start flats at J. Guide probe was not parked. Start flats at H. Guide probe was parked during these and for some the shadow is seen, for some not. Flats at Ks are good. These cover a range of ~5000 - 20000 counts.

-- OlgaKuhn - 12 Mar 2012
Topic revision: r5 - 12 Mar 2012, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback